W2

1. SAT scores are distributed with a mean of 1,500 and a standard deviation of 300. You are interested in estimating the average SAT score of first year students at your college. If you would like to limit the margin of error of your 98% confidence interval to 40 points, at least how many students should you sample?


A. 131
B. 216
C. 217
D.306

7. If it’s relatively riskier to reject the null hypothesis when it might be true, should a smaller or a larger significance level be used?

A. smaller
B. larger
1
Expert's answer
2020-06-08T21:22:08-0400

1) D.306

error=+/-z*sigma/"\sqrt n"

using table, critical valuefor P(0.98)=2.33

40=2.33*300/"\sqrt n"

"\sqrt n" =17.475

n=306

7) A. smaller

Lower values of α make it harder to reject the null hypothesis